# PAC-Chernoff Bounds: Understanding Generalization in the Interpolation Regime

### v3

@@ 1 Introduction
In modern machine learning, model classes have such large capacity that optimizers virtually always retrieve a model interpolating the training data; that is, with null training error (Zhang et al., 2017). These models are called interpolators and it is well known that there are many within a large model class (Livni et al., 2014) and that some of them have a remarkably small generalization error (difference between “training error” and “expected or test error”) while others do not (Feldman and Zhang, 2020).
The machine learning community has made a great effort during the last years to understand why modern learning algorithms retrieve, most of the time, interpolators with a small generalization error (Nagarajan and Kolter, 2019c; Bartlett et al., 2020). Most of these attempts often focus on providing generalization bounds. These bounds provide an upper limit on the expected error, tying it to variables related to the training dataset and the model produced by a learning algorithm. Such bounds typically resemble the following form,
@@
Examples of bounds falling inside this scheme include Vapnik–Chervonenkis (VC) bounds (Vapnik and Chervonenkis, 2015; Bartlett et al., 2019), Rademacher bounds (Mohri et al., 2018), (Norm and margin)-based bounds (Nagarajan and Kolter, 2019b; Neyshabur et al., 2015) and Sharpness-based measures (Nagarajan and Kolter, 2019a; Keskar et al., 2016). Even PAC-Bayes with a fixed prior (McAllester, 1999) can be included here if we consider algorithms retrieving a distribution over the models.
However, many recent works are increasingly suggesting that these bounds are probably vacuous in the over-parameterized regime and, in consequence, unable to explain generalization in this setting. Zhang et al. (2017) was the first one to provide quite convincing empirical evidence of this in the context of deep neural networks. Nagarajan and Kolter (2019c) later provided a series of experiments and theoretical results in the same direction. These works also show that even in the case where the complexity measure $\mathcal{C}$ depends directly on the algorithm $\mathcal{A}$ rather than simply in its output $\mathcal{A}(D)$ (for example, assuming the algorithm always retrieve models whose parameters’ norm is lower than a given constant) are also vacuous. Recently, Wang et al. (2024) showed how any near-interpolator exhibits rapid parameter norm growth, which implies that existing data-dependent parameter-norm-based bounds are necessarily loose, concluding that “explaining the generalization capability of near-interpolators will require new tools”. More generally, Gastpar et al. (2023) managed to show that, under some conditions resembling over-parameterization, there is no generalization bound like Equation (1) tight for all data-generating distributions, and that algorithm-dependent bounds are also provably vacuous. The emerging conclusion is that:
@@
These works (Nagarajan and Kolter, 2019c; Nagarajan, 2021; Gastpar et al., 2023) directly or indirectly advocate for the need of exploring alternative bounds that not solely depend on the training data but also use information of the data-generating distribution. These bounds are usually referred to as distribution-dependent bounds; where the complexity term ${\cal C}$ explicitly depends on the data-generating distribution $\nu$. Distribution-dependent bounds have been used before in different contexts (Zhang, 2006; Catoni, 2007), however, to the best of our knowledge, none of these bounds is known to be tight or, at least, non-vacuous for over-parameterized interpolators. Even more importantly, these kind of bounds have not been used before to analyze the generalization of over-parametrized interpolators.
@@ 1.1 Our Contribution
which ensures that the proposed bound is perfectly tight for interpolators, that is to say, to models whose empirical loss is null or very small, denoted as $\hat{L}({\cal A}(D),D)\leq\epsilon$.
As we show in this work, these bounds are directly connected to Large Deviation Theory (LDT) (Ellis, 2012) because their complexity measure ${\cal C}({\cal A}(D),n,\nu,\delta)$ directly depends on the so-called rate function (also known as the Cramér-Chernoff function), which is the central element of LDT. We employ the rate function to present a new characterization of the smoothness of a model using distribution-dependent measures. According to Theorem 4.5, this approach enables a precise characterization of which interpolators better generalize, addressing an outstanding open question in machine learning.
@@
We do not claim that the proposed theoretical analyses of each of these learning techniques are better than existing ones at individual level. For example, we do not claim that our explanation about why $\ell_{2}$-norm promotes generalization is better than existing ones, for example, in the context of linear regression (Bartlett et al., 2020). However, we do claim that using the rate function and PAC-Chernoff bounds, which are distribution-dependent quantities, it is possible to jointly explain, up to some degree, the generalization of interpolators, the role of over-parameterization and why many widely used modern learning techniques find interpolators that generalize. To the best of our knowledge, this is the first theoretical framework capable of achieving insights that covers such a wide range of learning techniques. At the same time, this work shows that distribution-dependent bounds are a promising research direction for understanding the generalization of modern machine learning methods.
@@ 2 Preliminaries
First of all, let us introduce the notation and the assumptions of this work. Let $D$ denote a given training data set with size $n>0$, which we assume has been i.i.d. generated from an unknown distribution, denoted by $\nu({\boldsymbol{y}},{\boldsymbol{x}})$. We assume we have a model class parameterized by a parameter vector ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$. We denote for any model ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$, its loss as $\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$, its expected log-loss as $L({\boldsymbol{\theta}})={\mathbb{E}}_{\nu}[\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})]$, and its empirical log-loss as $\hat{L}({\boldsymbol{\theta}},D)=\frac{1}{n}\sum_{i}\ell({\boldsymbol{y}}_{i},{\boldsymbol{x}}_{i},{\boldsymbol{\theta}})$. Finally, the subset of models having a null variance (constant predictions) is denoted by ${\boldsymbol{\Theta}}_{0}=\{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}:\mathbb{V}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}}))=0\}$.
We also consider that our model class lives inside a finite-precision computing machine. In that sense, we assume our model class is represented by $p$ parameters. Then, we will have $k^{p}$ different models, under a numeric precision of $\log_{2}(k)$ bits (for example, the value of $k$ would be $k=2^{32}$ for single floating point precision and $k=2^{64}$ for double floating point precision). This model class defines a finite grid (using a discretization approach) contained in $\mathbb{R}^{p}$, but it does not imposed any restriction in the definition of the loss, which can be defined in a continuous space. In Section 8, we discuss how this restriction could be potentially lifted by using recently introduced PAC-Bayes-Chernoff bounds (Casado et al., 2024). Let us now introduce the specific assumptions of this work:
@@ 3 The Rate function
We start introducing the key function of this work, the so-called rate function, denoted by ${\cal I}_{\boldsymbol{\theta}}(a)$. The rate function places a central role in Large Deviation Theory (LDT) (Ellis, 2012), a branch of probability theory tightly connected to statistical mechanics, that deals with understanding the behavior of rare events or large fluctuations in random systems. The rate function is normally used to understand the underlying structure of these events and how their probabilities change as they move away from the typical or average behavior.
@@
Note that, when $\mathbb{P}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})=m_{\boldsymbol{\theta}})\neq 0$, for $\forall s\geq-\ln\mathbb{P}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})=m_{\boldsymbol{\theta}})$, ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(s\right)$ is constantly equal to $L({\boldsymbol{\theta}})-m_{\boldsymbol{\theta}}$ and, in consequence, ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(s\right)$ is a generalized inverse of ${\cal I}_{\boldsymbol{\theta}}(a)$ (Rockafellar, 1970). For the sake of simplicity, we note that we abuse notation and assume throughout the rest of work that for $a>L({\boldsymbol{\theta}})-m_{\boldsymbol{\theta}}$, ${\cal I}_{\boldsymbol{\theta}}(a)=\infty$. We also assume that binary operators apply to this value following common sense. According to the following proposition, both the rate function and its inverse are well defined for models satisfying Assumption 1.
@@
The relevance of the rate function is consequence of the following results; the first one, the classic Chernoff bound, defines how likely is to observe, over different i.i.d. data sets, an empirical loss $\hat{L}(D,{\boldsymbol{\theta}})$ that deviates from the expected loss $L({\boldsymbol{\theta}})$ by a positive quantity $a>0$.
Theorem 3.3 (Chernoff (1952)).
@@
The above bound is relevant in this work because of two main factors that characterize its behavior w.r.t. the data set size $n$ and the generalization error gap $a$. On one hand, Chernoff’s bound is known to be quite loose in the mean of the variable (when $(\text{when }a\approx 0)$ but tight on the tail (when $(\text{when }a\approx L({\boldsymbol{\theta}})-m_{{\boldsymbol{\theta}}})$. As a result, the bound is specially useful when talking about interpolators, where the value of $L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}})$ is close to its maximum possible value $(a=L({\boldsymbol{\theta}})-m_{{\boldsymbol{\theta}}})$, as stated in the following result.
@@
On the other hand, Cramér’s Theorem (Cramér, 1938) states that Chernoff’s bound is exponentially tight for large $n$. Formally, this statement is written as follows,
Theorem 3.5 (Cramér (1938); Ellis (2012)).
@@
which shows that the exact expression of ($\mathbb{P}{}_{D\sim\nu^{n}}(L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}})\geq a)$ is defined by the rate function, up to a sub-exponential term that can be discarded when $n$ is large, because $\lim_{n\rightarrow\infty}\tfrac{o(n,a)}{n}=0$ and, in consequence, it does not have any meaningful effect. Then, according to LDT, when $n$ is large, the rate function would be the key quantity describing the generalization error of a model, that is, statistical behavior of the difference between the expected $L({\boldsymbol{\theta}})$ and the empirical loss $\hat{L}(D,{\boldsymbol{\theta}})$.
@@
The Chernoff bound states that the generalization error a model is governed by its rate function. The rate function is an oracle or distribution-dependent quantity because it depends on the data-generating distribution $\nu$. In machine learning, this distribution $\nu$ is always unknown, which may be seen as a problem. However, as discussed in the introduction, we will show how the use of a distribution-dependent quantity is able to explain, for example, why data-augmentation works, why we need over-parameterized model families, and why we need invariant architectures, like convolutional neural networks. Furthermore, in this work, we also show how the rate function can be easily estimated from an independent data set, if it is used as a proxy for the data-generating distribution. Something commonly done in machine learning to get estimates of $L({\boldsymbol{\theta}})$. In Appendix A, we show that the rate function ${\cal I}_{\boldsymbol{\theta}}(\cdot)$ can be easily estimated by running the model once on an independent data set, using log-sum-exp operations and a grid search. The fact that this relevant function can be easily estimated and plotted is, in our opinion, a relevant finding of this work. This will allow us to empirically illustrate and validate the predictions made by our theoretical framework.
@@
Let us briefly illustrate why the rate function can help to understand the generalization of interpolators. Figure 2 displays an estimation of the rate function ${\cal I}_{\boldsymbol{\theta}}(\cdot)$ for some neural network models used in Zhang et al. (2017). This work demonstrated that, within the same model class, interpolators that merely memorize the training data (as represented by “Random” in the figure, which has been learned using a random-labeled training data set) can co-exist with others that generalize exceptionally well, as the ones learned using weight-norm regularization (labeled as “L2”) and/or data-augmentation techniques (random-cropping labeled as “Crop”). According to Chernoff’s bound, models with larger rate functions are less likely to have significant disparities between their expected and empirical losses or, in other words, the empirical loss $\hat{L}(D,{\boldsymbol{\theta}})$ is more concentrated around its mean $L({\boldsymbol{\theta}})$. Figure 3 illustrates this fact for three of these models by plotting histograms that showcase the distribution of $\hat{L}(D,{\boldsymbol{\theta}})$ across various data sets of size $n=50$ (retrieved from the test-set). From the histograms, it is evident that the concentration of $\hat{L}(D,{\boldsymbol{\theta}})$ varies among the models; the initial model, defined by Kaiming or He initialization (Goodfellow et al., 2016), has a prominent rate function. For this model, $\hat{L}(D,{\boldsymbol{\theta}})$ is tightly concentrated around its mean, $L({\boldsymbol{\theta}})=\ln 10$, observable through the minute scale of its x-axis. Comparatively, the Standard model, trained using SGD and characterized by a smaller rate function, has a more dispersed distribution around its mean $L({\boldsymbol{\theta}})=0.65$. This dispersion is notably wider than that of the L2-Crop model. This latter model, also trained with SGD but incorporating both $\ell_{2}$ regularization and data-augmentation, has a larger rate function than the Standard model.
@@
The main point of Figures 2 and 3 is to illustrate how $\hat{L}(D,{\boldsymbol{\theta}})$ is a random variable that could be either highly or poorly concentrated around its mean $L({\boldsymbol{\theta}})$, something which is captured by the rate function of the model. When a model interpolates the training data, we must think we are observing a realization of this random variable $\hat{L}(D,{\boldsymbol{\theta}})$ for a particular data set $D$. Consequently, when comparing two models that interpolate the training data, we should choose the one that is more concentrated around its mean (that is, with a higher rate function); because it will likely have a smaller expected loss. For example, Figure 3 (left) and (center) describe the distribution of $\hat{L}(D,{\boldsymbol{\theta}})$ for two models, Standard and L2-Crop, respectively. As both models interpolate the training data, $\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon$, it means we should prefer the L2-Crop model because $\hat{L}(D,{\boldsymbol{\theta}})$ is more concentrated around its expected value (its rate function is also higher). In fact, as shown in Figure 2, the L2-Crop model is the one generalizing better.

### v4

@@ 1 Introduction
In modern machine learning, model classes have such large capacity that optimizers virtually always retrieve a model interpolating the training data; that is, with null training error (Zhang et al.,, 2017). These models are called interpolators and it is well known that there are many within a large model class (Livni et al.,, 2014) and that some of them have a remarkably small generalization error (difference between “training error” and “expected or test error”) while others do not (Feldman & Zhang,, 2020).
The machine learning community has made a great effort during the last years to understand why modern learning algorithms retrieve, most of the time, interpolators with a small generalization error (Nagarajan & Kolter, 2019b, ; Bartlett et al.,, 2020). Most of these attempts often focus on providing generalization bounds. These bounds provide an upper limit on the expected error, tying it to variables related to the training dataset and the model produced by a learning algorithm. Such bounds typically resemble the following form,
@@
Examples of bounds falling inside this scheme include Vapnik–Chervonenkis (VC) bounds (Vapnik & Chervonenkis,, 2015; Bartlett et al.,, 2019), Rademacher bounds (Mohri et al.,, 2018), (Norm and margin)-based bounds (Nagarajan & Kolter,, 2017; Neyshabur et al.,, 2015) and Sharpness-based measures (Nagarajan & Kolter, 2019a, ; Keskar et al., 2017a, ). Even PAC-Bayes with a fixed prior (McAllester,, 1999) can be included here if we consider algorithms retrieving a distribution over the models.
However, many recent works are increasingly suggesting that these bounds are probably vacuous in the over-parameterized regime and, in consequence, unable to explain generalization in this setting. Zhang et al., (2017) was the first one to provide quite convincing empirical evidence of this in the context of deep neural networks. Nagarajan and Kolter (2019b) later provided a series of experiments and theoretical results in the same direction. These works also show that even in the case where the complexity measure $\mathcal{C}$ depends directly on the algorithm $\mathcal{A}$ rather than simply in its output $\mathcal{A}(D)$ (for example, assuming the algorithm always retrieve models whose parameters’ norm is lower than a given constant) are also vacuous. Recently, Wang et al., (2024) showed how any near-interpolator exhibits rapid parameter norm growth, which implies that existing data-dependent parameter-norm-based bounds are necessarily loose, concluding that “explaining the generalization capability of near-interpolators will require new tools”. More generally, Gastpar et al., (2024) managed to show that, under some conditions resembling over-parameterization, there is no generalization bound like Equation (1) tight for all data-generating distributions, and that algorithm-dependent bounds are also provably vacuous. For all these reasons, the emerging conclusion is that:
@@
The aforementioned works (Nagarajan & Kolter, 2019b, ; Nagarajan,, 2021; Gastpar et al.,, 2024) directly or indirectly advocate for the need of exploring alternative bounds that not solely depend on the training data but also use information of the data-generating distribution. These bounds are usually referred to as distribution-dependent bounds; where the complexity term ${\cal C}$ explicitly depends on the data-generating distribution $\nu$. Distribution-dependent bounds have been used before in different contexts (Zhang,, 2006; Catoni,, 2007), however, to the best of our knowledge, none of these bounds is known to be tight or, at least, non-vacuous for over-parameterized interpolators. Even more importantly, these kind of bounds have not been used before to analyze the generalization of over-parametrized interpolators.
@@ 1.1 Our Contribution
which ensures that the proposed bound is perfectly tight for interpolators, that is to say, to models whose empirical loss is null or small enough to be considered negligible or within an acceptable error margin, denoted as $\hat{L}({\cal A}(D),D)\leq\epsilon$.
As we show in this work, these bounds are directly connected to Large Deviation Theory (LDT) (Ellis,, 2012) because their complexity measure ${\cal C}({\cal A}(D),n,\nu,\delta)$ directly depends on the so-called rate function (also known as the Cramér-Chernoff function), which is the central element of LDT. We employ the rate function to present a new characterization of the smoothness of a model using distribution-dependent measures. According to Theorem 4.6, this approach enables a precise characterization of which interpolators better generalize, addressing an outstanding open question in machine learning.
@@
We do not claim that the proposed theoretical analyses of each of these learning techniques are better than existing ones at individual level. For example, we do not claim that our explanation about why $\ell_{2}$-norm promotes generalization is better than existing ones, for example, in the context of linear regression (Bartlett et al.,, 2020). However, we do claim that using the rate function and PAC-Chernoff bounds, which are distribution-dependent quantities, it is possible to jointly explain, up to some degree, the generalization of interpolators, the role of over-parameterization and why many widely used modern learning techniques find interpolators that generalize. To the best of our knowledge, this is the first theoretical framework capable of achieving insights that covers such a wide range of learning techniques. At the same time, this work shows that distribution-dependent bounds are a promising research direction for understanding the generalization of modern machine learning methods.
@@ 2 Preliminaries
First of all, let us introduce the notation and the assumptions of this work. Let $D$ denote a given training data set with size $n>0$, which we assume has been i.i.d. generated from an unknown distribution, denoted by $\nu({\boldsymbol{y}},{\boldsymbol{x}})$. We assume we have a model class parameterized by a parameter vector ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$. We denote for any model ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$, its loss as $\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$, its expected loss as $L({\boldsymbol{\theta}})={\mathbb{E}}_{\nu}[\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})]$, and its empirical loss as $\hat{L}({\boldsymbol{\theta}},D)=\frac{1}{n}\sum_{i}\ell({\boldsymbol{y}}_{i},{\boldsymbol{x}}_{i},{\boldsymbol{\theta}})$. Finally, the subset of models having a null variance (constant predictions) is defined as ${\boldsymbol{\Theta}}_{0}:=\{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}:\mathbb{V}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}}))=0\}$.
We also consider that our model class lives inside a finite-precision computing machine. In that sense, we assume our model class is represented by $p$ parameters. Then, we will have $k^{p}$ different models, under a numeric precision of $\log_{2}(k)$ bits (for example, the value of $k$ would be $k=2^{32}$ for single floating point precision and $k=2^{64}$ for double floating point precision). This model class defines a finite grid (using a discretization approach) contained in $\mathbb{R}^{p}$, but it does not imposed any restriction in the definition of the loss, which can be defined in a continuous space. In Section 10, we discuss how the results presented here could be potentially extended to infinite model classes by using recently introduced PAC-Bayes-Chernoff bounds (Casado et al.,, 2024). Let us now introduce the specific assumptions of this work:
@@ 3 The Rate Function
We start introducing the key function of this work, the so-called rate function, denoted by ${\cal I}_{\boldsymbol{\theta}}(a)$. The rate function places a central role in Large Deviation Theory (LDT) (Ellis,, 2012), a branch of probability theory tightly connected to statistical mechanics, that deals with understanding the behavior of rare events or large fluctuations in random systems. The rate function is normally used to understand the underlying structure of these events and how their probabilities change as they move away from the typical or average behavior.
@@
Note that, when $\mathbb{P}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})=m_{\boldsymbol{\theta}})\neq 0$, for $\forall s\geq-\ln\mathbb{P}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})=m_{\boldsymbol{\theta}})$, ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(s\right)$ is constantly equal to $L({\boldsymbol{\theta}})-m_{\boldsymbol{\theta}}$ and, in consequence, ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(s\right)$ is a generalized inverse of ${\cal I}_{\boldsymbol{\theta}}(a)$ (Rockafellar,, 1970). For the sake of simplicity, we note that we abuse notation and assume throughout the rest of work that for $a>L({\boldsymbol{\theta}})-m_{\boldsymbol{\theta}}$, ${\cal I}_{\boldsymbol{\theta}}(a)=\infty$. We also assume that binary operators apply to this value following common sense. According to the following proposition, both the rate function and its inverse are well defined for models satisfying Assumption 1.
@@
The relevance of the rate function is consequence of the following results; firstly, the classic Chernoff bound, defines how likely is to observe, over different i.i.d. data sets, an empirical loss $\hat{L}(D,{\boldsymbol{\theta}})$ that deviates from the expected loss $L({\boldsymbol{\theta}})$ by a positive quantity $a>0$.
Theorem 3.3 (Chernoff, (1952)).
@@
The above bound is relevant in this work because of two main factors that characterize its behavior w.r.t. the data set size $n$ and the generalization error gap $a$: on one hand, Chernoff’s bound is known to be quite loose in the mean of the variable (when $(\text{when }a\approx 0)$ but tight on the tail (when $(\text{when }a\approx L({\boldsymbol{\theta}})-m_{{\boldsymbol{\theta}}})$. As a result, the bound is specially useful when talking about interpolators, where the value of $L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}})$ is close to its maximum possible value $(a=L({\boldsymbol{\theta}})-m_{{\boldsymbol{\theta}}})$, as stated in the following result.
@@
On the other hand, Cramér’s Theorem (Cramér,, 1938) states that Chernoff’s bound is exponentially tight for large $n$. Formally, this statement is written as follows,
Theorem 3.5 (Cramér, (1938); Ellis, (2012)).
@@
which shows that the exact expression of ($\mathbb{P}{}_{D\sim\nu^{n}}(L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}})\geq a)$ is defined by the rate function, up to a sub-exponential term, that is negligible when $n$ is large, because $\lim_{n\rightarrow\infty}\tfrac{o(n,a)}{n}=0$ and, in consequence, it does not have any meaningful effect. Then, according to LDT, when $n$ is large, the rate function would be the key quantity describing the generalization error of a model, that is, statistical behavior of the difference between the expected $L({\boldsymbol{\theta}})$ and the empirical loss $\hat{L}(D,{\boldsymbol{\theta}})$.
@@
The Chernoff bound states that the generalization error a model is governed by its rate function. The rate function is an oracle and distribution-dependent quantity because it depends on the data-generating distribution $\nu$. In machine learning, this distribution $\nu$ is always unknown, which may be considered as a problem. However, as discussed in the introduction, we will show how the use of a distribution-dependent quantity allows to explain, for example, why data-augmentation works, why we need over-parameterized model families, and why we should consider invariant architectures, like convolutional neural networks. Furthermore, in this work, we also show how the rate function can be easily estimated from an independent data set, if it is used as a proxy for the data-generating distribution. This is a common practice in machine learning by making use of a separated validation data set to get estimates of $L({\boldsymbol{\theta}})$ without incurring in data snooping. In Appendix A, we show that the rate function ${\cal I}_{\boldsymbol{\theta}}(\cdot)$ can be easily estimated by running the model once on this validation data set and using log-sum-exp operations to estimate $J_{{\boldsymbol{\theta}}}(\lambda)$; and using a simple grid search to optimize the value of $\lambda$ in Definition 1. The fact that the cumulant and rate function can be easily estimated and plotted (as in Figure 2) is, in our opinion, a relevant finding of this work. This will allow us to empirically illustrate and validate the predictions made by our theoretical framework.
@@
Let us briefly illustrate why the rate function is useful to understand the generalization of interpolators. Figure 2 displays an estimation of the rate function ${\cal I}_{\boldsymbol{\theta}}(\cdot)$ for some neural network models used in Zhang et al., (2017). This work demonstrated that, within the same model class, interpolators that merely memorize the training data (as represented by “Random” in the figure, which has been learned using a random-labeled training data set) can co-exist with others that generalize exceptionally well, as the ones learned using weight-norm regularization (labeled as “L2”) and/or data-augmentation techniques (random-cropping labeled as “Crop”). According to Chernoff’s bound, models with a larger rate function are less likely to have significant disparities between their expected and empirical losses; in other words, the empirical loss $\hat{L}(D,{\boldsymbol{\theta}})$ is more concentrated around its mean $L({\boldsymbol{\theta}})$. Figure 3 illustrates this fact for three of these models by plotting histograms that showcase the distribution of $\hat{L}(D,{\boldsymbol{\theta}})$ across various data sets of size $n=50$ (retrieved from the test set). From the histograms, it is clear that the concentration of $\hat{L}(D,{\boldsymbol{\theta}})$ varies among the models; the initial model, defined by Kaiming or He initialization (Goodfellow et al.,, 2016), has a prominent rate function. For this model, $\hat{L}(D,{\boldsymbol{\theta}})$ is tightly concentrated around its mean, $L({\boldsymbol{\theta}})=\ln 10$, observable through the minute scale of its x-axis. Comparatively, the Standard model, trained using SGD and characterized by a smaller rate function, has a more dispersed distribution around its mean $L({\boldsymbol{\theta}})=0.65$. This dispersion is notably wider than that of the L2-Crop model. The latter, also trained with SGD but incorporating both $\ell_{2}$ regularization and data-augmentation, has a larger rate function than the Standard model.
@@
The main point of Figures 2 and 3 is to illustrate how $\hat{L}(D,{\boldsymbol{\theta}})$ is a random variable that could be either highly or poorly concentrated around its mean $L({\boldsymbol{\theta}})$, something which is captured by the rate function of the model. When a model interpolates the training data, we must think we are observing a realization of this random variable $\hat{L}(D,{\boldsymbol{\theta}})$ for a particular data set $D$. Consequently, when comparing two models that interpolate the training data, we should choose the one that is more concentrated around its mean (higher rate function); because it will likely have a smaller expected loss. For example, Figure 3 (left) and (center) describe the distribution of $\hat{L}(D,{\boldsymbol{\theta}})$ for two models, Standard and L2-Crop, respectively. As both models interpolate the training data, we should prefer L2-Crop because $\hat{L}(D,{\boldsymbol{\theta}})$ is more concentrated around its expected value for that model.